OpenStudy (anonymous):

HELP PLEASE Find the standard form of the equation of the parabola with a focus at (0, -3) and a directrix at y = 3.

OpenStudy (anonymous):

all points such that distance to focus = distance to directrix. (y-3)^2 = (x-0)^2 + (y+3)^2 y=(x-3)^2 / 6 +9/6

OpenStudy (anonymous):

Thank you